OpenStudy (anonymous):

Mary wants to hang a mirror in her room but the mirror and frame must not have an area larger than 7 square feet. The mirror is 2 feet wide and 3 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The frame must have equal width of x on each side.)

OpenStudy (anonymous):

x2 + 14x − 2 = 0 2x2 + 10x − 7 = 0 3x2 + 12x − 7 = 0 4x2 + 10x − 1 = 0